OpenStudy (anonymous):

Find the domain of the given function. f(x) = square root of quantity x plus three divided by quantity x plus eight times quantity x minus two.

OpenStudy (anonymous):

Square of quantity \(x\) plus three means : \(\sqrt{x+3}\) quantity \(x\) plus eight means : \((x+8)\) quantity \(x\) minus two means : \((x - 2)\) So I interpreted it as : \[f(x) = \frac{\sqrt{x + 3}}{(x + 8)} \times (x-2)\]

OpenStudy (anonymous):

See, in square, whatever the thing is, it cannot become Negative at any point of time. So: \[(x+3) \ge 0 \implies x \ge -3\] This is first condition..

OpenStudy (anonymous):

Second condition is, the thing in denominator cannot become \(0\).. \[x+8 \ne 0 \implies x \ne -8\]

OpenStudy (anonymous):

So, Domain is : All Reals greater than or equal 3..

OpenStudy (anonymous):

*-3..

OpenStudy (anonymous):

\[\mathbb{R} : x \ge -3\]